OpenStudy (anonymous):

What is the quotient of t^2-16/6t ÷2t+8/t-4 ?

OpenStudy (anonymous):

\[ \frac{t^2-16}{6t} ÷\frac{2t+8}{t-4}\\ \frac{t^2-16}{6t} \frac{t-4}{2t+8}\\ \frac{(t -4)(t+4)}{6t} \frac{t-4}{2(t+4)}\\ \] Can you finish it from here?

OpenStudy (anonymous):

i got t^2-8t+16/12t but i know it is either a or b: A) t^2-8t+16/12t, t ≠ −4, 0, 4 B) t^2-8t+16/12t, t ≠ 0, 4 which do you think?

OpenStudy (anonymous):

I will say A. since the the fraction or its variants are not defined for t=-4,0,4

OpenStudy (anonymous):

okay, thank you